find the class boundaries, midpoint, and width for the class 12.4-15.2

1
Expert's answer
2022-05-23T13:41:14-0400

Lower Boundary: 12.4-0.05=12.35.

Upper Boundary: 15.2+0.05=15.25.


Midpoint: "\frac{12.4+15.2}{2}=13.8."


Width: "15.25-12.35=2.9"